# Background Semantics Matter: Cross-Task Feature Exchange Network for Clustered Infrared Small Target Detection With Sky-Annotated Dataset

### v1

@@ V-A Experimental Settings
TABLE III: Quantitative ablation study evaluates the impact of the Background-Aware segmentation, Feature Exchange module, and adapter module within the Background-Aware Feature Exchange Network (BAFE-Net)
@@
TABLE IV: Quantitative ablation study was conducted to evaluate the impact of the number of hidden layers in the Adapter
@@
TABLE V: Quantitative ablation study evaluates the impact of different Feature Exchange Mechanisms within the Feature Exchange module of BAFE-Net
@@
TABLE VI: Quantitative ablation study evaluates the impact of different backbones on BAFE-Net performance
@@
  (a) ResNet18 0.258 0.328 0.233 0.341
(b) ResNet34 0.258 0.330 0.226 0.332
(c) ResNet50 0.270 0.332 0.236 0.329
(d) ResNet101 0.256 0.324 0.241 0.336
TABLE VII: Comparison with Other State-of-the-art methods on DenseSIRST.
Method Backbone DenseSIRST
mAP07 $\uparrow$ recall07 $\uparrow$ mAP12 $\uparrow$ recall12 $\uparrow$ Flops $\downarrow$ Params $\downarrow$
              One-stage
FCOS [34] ResNet50 0.232 0.315 0.204 0.324 50.291G 32.113M
SSD [36] 0.211 0.421 0.178 0.424 87.552G 23.746M
GFL [37] ResNet50 0.253 0.332 0.230 0.317 52.296G 32.258M
ATSS [38] ResNet50 0.248 0.327 0.202 0.326 51.504G 32.113M
CenterNet [39] ResNet50 0.000 0.000 0.000 0.000 50.278G 32.111M
PAA [40] ResNet50 0.255 0.545 0.228 0.551 51.504G 32.113M
PVT-T [41] 0.109 0.481 0.093 0.501 41.623G 21.325M
RetinaNet [42] ResNet50 0.114 0.510 0.086 0.523 52.203G 36.330M
EfficientDet [43] 0.099 0.433 0.072 0.419 34.686G 18.320M
TOOD [13] ResNet50 0.256 0.355 0.226 0.342 50.456G 32.018M
VFNet [44] ResNet50 0.253 0.336 0.214 0.336 48.317G 32.709M
YOLOF [45] ResNet50 0.091 0.009 0.002 0.009 25.076G 42.339M
AutoAssign [46] ResNet50 0.255 0.354 0.180 0.314 50.555G 36.244M
DyHead [47] ResNet50 0.249 0.335 0.189 0.328 27.866G 38.890M
Two-stage
Faster R-CNN [48] ResNet50 0.091 0.022 0.015 0.029 0.759T 33.035M
Cascade R-CNN [49] ResNet50 0.136 0.188 0.139 0.194 90.978G 69.152M
Dynamic R-CNN [50] ResNet50 0.184 0.235 0.111 0.190 63.179G 41.348M
Grid R-CNN [51] ResNet50 0.091 0.018 0.025 0.037 0.177T 64.467M
Libra R-CNN [52] ResNet50 0.141 0.142 0.085 0.120 63.990G 41.611M
End2End
DETR [53] ResNet50 0.000 0.000 0.000 0.000 24.940G 41.555M
Deformable DETR [54] ResNet50 0.024 0.016 0.018 0.197 51.772G 40.099M
DAB-DETR [55] ResNet50 0.005 0.054 0.000 0.001 28.939G 43.702M
Conditional DETR [56] ResNet50 0.000 0.000 0.000 0.001 27.143G 40.297M
Sparse R-CNN [57] ResNet50 0.183 0.572 0.154 0.614 45.274G 0.106G
$\star$ BAFE-Net (Ours) ResNet50 0.270 0.332 0.236 0.329 69.114G 35.329M
V-B Ablation Study
V-B1 Ablation Study on Background-Aware Gaussian Copy-Paste
In this study, we initiate our investigation from the data augmentation module, focusing on the progressive integration of our proposed Copy-Paste technique, clustering of pasted objects, Gaussian module, and Background-Aware pasting. Our aim is to empirically validate the enhancements to detection performance brought about by these steps. To achieve this, we explore various combinations of the aforementioned components and conduct experimental validation of the results.
The experimental outcomes, as outlined in Table II, generate several key insights, providing a holistic understanding of the impact of each module. Firstly, the effective enhancement of detection performance by each module is noticeable, corroborating the initial motivation of our study. The Copy-Paste technique, even when used independently, significantly boosts performance, as evidenced by the increases of 1.1% and 2.3% on the DenseSIRST dataset, when comparing strategy (a) with strategy (b).
V-B2 Ablation Study on the Scheme of Background-Aware Feature Exchange
The Background-Aware Feature Exchange module constitutes the core contribution of this paper, as evidenced by its significant impact on detection performance demonstrated in Table III. To validate the effectiveness of the proposed components, we conducted an extensive ablation study by exploring various combinations of the background-based segmentation, Feature Exchange module, and adapter module.
Table III encapsulates the results of our experimental study and yields several crucial insights. Firstly, each constituent module substantively and independently enhances the detection performance, thereby validating the underpinnings of our research. Secondly, comparison between Strategy (a) and Strategy (b) reveals a significant performance augmentation when background-based segmentation is incorporated into object detection, wherein the detection head interacts with the segmentation head. This integration accounted for an increase of 1.0% and 0.3% on the DenseSIRST dataset.
As the feature adapter employs a Multi-Layer Perceptron (MLP), the number of hidden layers in the MLP is a crucial parameter that warrants careful consideration. Increasing the number of hidden layers typically improves performance but comes at the cost of increased computational complexity. To strike a balance between performance and efficiency, we conducted an experiment to determine the optimal number of hidden layers, as shown in Table  IV. The experimental results demonstrate that increasing the number of hidden layers leads to a consistent improvement in detection performance. However, the performance gain becomes less significant when the number of hidden layers exceeds 64, while the parameter count increases substantially. Therefore, we opted for 64 hidden layers in the Adapter, as it provides a significant performance boost while maintaining a reasonable computational overhead compared to using 128 hidden layers or more.
V-B3 Ablation Study on Feature Exchange Mechanism
As delineated in Table V, the superiority of the Channel Exchange method over alternative Feature Exchange mechanisms is clearly manifested. The empirical evidence underscores the robustness and optimality of Channel Exchange, thereby validating its adoption as the preferred method for Feature Exchange in our model.
V-B4 Impact of Backbone Depth:
We conducted a series of experiments using different ResNet architectures, varying the depth from ResNet18 to ResNet101, to identify the optimal backbone for our proposed BAFE-Net. The results of these experiments are summarized in Table VI. Notably, ResNet50 emerged as the most suitable backbone for the BAFE-Net.
V-C Comparison with State-of-the-Arts
We benchmarked our BAFE-Net against current leading or state-of-the-art methods on the proposed DenseSIRST dataset, using metrics such as $mAP_{07}$, $mAP_{12}$, and Recall. The experimental results are presented in Table VII.
V-D Visual Analysis

### v2

@@ V-B Comparison with State-of-the-Arts
TABLE IV: Quantitative ablation study evaluates the impact of the Background-Aware segmentation, Feature Exchange module, and adapter module within the Background-Aware Feature Exchange Network (BAFE-Net)
@@
TABLE V: Quantitative ablation study was conducted to evaluate the impact of the number of hidden layers in the Adapter
@@
TABLE VI: Quantitative ablation study evaluates the impact of different Feature Exchange Mechanisms within the Feature Exchange module of BAFE-Net
@@
TABLE VII: Ablation study on the impact of feature selection on detection performance in Channel-wise Feature Exchange
Strategy Feature Selection DenseSIRST
mAP07 $\uparrow$ recall07 $\uparrow$ mAP12 $\uparrow$ recall12 $\uparrow$
  (a) Random Selection 0.255 0.322 0.172 0.317
(b) Fixed Selection 0.273 0.333 0.230 0.323
(c) Dynamic Selection 0.283 0.335 0.233 0.325
TABLE VIII: Ablation study on the impact of the number of exchange channels on detection performance in Top-K selection
Strategy P DenseSIRST
mAP07 $\uparrow$ recall07 $\uparrow$ mAP12 $\uparrow$ recall12 $\uparrow$
  (a) 1/2 0.283 0.335 0.233 0.325
(b) 1/4 0.241 0.316 0.230 0.331
(c) 3/4 0.242 0.325 0.204 0.310
(d) 1/8 0.238 0.312 0.218 0.323
(e) 3/8 0.255 0.327 0.203 0.331
(f) 5/8 0.259 0.325 0.201 0.328
(g) 7/8 0.238 0.316 0.214 0.329
(h) 1/16 0.245 0.326 0.213 0.301
(i) 1/32 0.236 0.319 0.212 0.312
TABLE IX: Quantitative ablation study evaluates the impact of different backbones on BAFE-Net performance
@@
  (a) ResNet18 0.283 0.335 0.233 0.341
(b) ResNet34 0.281 0.354 0.226 0.332
(c) ResNet50 0.274 0.342 0.248 0.338
(d) ResNet101 0.261 0.317 0.241 0.336
V-C Visual Analysis
@@
V-D Ablation Study
V-D1 Ablation Study on Background-Aware Gaussian Copy-Paste
We initiate our investigation from the data augmentation module, focusing on the progressive integration of our proposed Copy-Paste technique, clustering of pasted objects, Gaussian module, and Background-Aware pasting. The experimental outcomes, as outlined in Table III, generate several key insights, providing a holistic understanding of the impact of each module. Firstly, the effective enhancement of detection performance by each module is noticeable, corroborating the initial motivation of our study. The Copy-Paste technique, even when used independently, significantly boosts performance, as evidenced by the increases of 1.1% and 2.3% on the DenseSIRST dataset, when comparing strategy (a) with strategy (b).
V-D2 Ablation Study on the Scheme of Background-Aware Feature Exchange
The Background-Aware Feature Exchange module constitutes the core contribution of this paper, as evidenced by its significant impact on detection performance demonstrated in Table IV. To validate the effectiveness of the proposed components, we conducted an extensive ablation study by exploring various combinations of the background-based segmentation, Feature Exchange module, and adapter module.
Table IV encapsulates the results of our experimental study and yields several crucial insights. Firstly, each constituent module substantively and independently enhances the detection performance, thereby validating the underpinnings of our research. Secondly, comparison between Strategy (a) and Strategy (b) reveals a significant performance augmentation when background-based segmentation is incorporated into object detection, wherein the detection head interacts with the segmentation head. This integration accounted for an increase of 1.0% and 0.3% on the DenseSIRST dataset.
As the feature adapter employs a Multi-Layer Perceptron (MLP), the number of hidden layers in the MLP is a crucial parameter that warrants careful consideration. Increasing the number of hidden layers typically improves performance but comes at the cost of increased computational complexity. To strike a balance between performance and efficiency, we conducted an experiment to determine the optimal number of hidden layers, as shown in Table  V. The experimental results demonstrate that increasing the number of hidden layers leads to a consistent improvement in detection performance. However, the performance gain becomes less significant when the number of hidden layers exceeds 64, while the parameter count increases substantially. Therefore, we opted for 64 hidden layers in the Adapter, as it provides a significant performance boost while maintaining a reasonable computational overhead compared to using 128 hidden layers or more.
V-D3 Ablation Study on Feature Exchange Mechanism
As delineated in Table VI, the superiority of the Channel Exchange method over alternative Feature Exchange mechanisms is clearly manifested. The empirical evidence underscores the robustness and optimality of Channel Exchange, thereby validating its adoption as the preferred method for Feature Exchange in our model.
V-D4 Ablation Study in Dynamic Channel Selection
In this section, we first conduct a rigorous exploration of feature selection methods. This includes random selection, fixed selection, and dynamic selection. The task is to determine the most effective feature selection method. 1) Randomly Selection: This method selects feature channels randomly, which may lead to suboptimal feature exchanges. In some cases, it could even select features that hinder the task, ultimately reducing model performance. 2) Fixed Selection: While fixed selection provides consistency during training and eliminates uncertainty, it lacks the adaptability to dynamic input features, thus limiting the model’s performance in complex or varied scenarios. 3) Dynamic Selection: dynamic selection maximizes the performance of the model and reduces unnecessary information redundancy by selecting the most useful feature channels for exchange. The exchange ratio is judiciously set to 0.5 to maintain a balance.
As shown in Table VII, the advantages of dynamic selection are obvious. Empirical evidence highlights the robustness and optimality of dynamic selection, thus confirming its adoption as the preferred method for feature selection in our model.
In addition, we conduct a rigorous exploration of the value of $p$ in Top-K selection. As shown in Table VIII, $p$ is optimal when it is 1/2.
V-D5 Impact of Backbone Depth
We conducted a series of experiments using different ResNet architectures, varying the depth from ResNet18 to ResNet101, to identify the optimal backbone for our proposed BAFE-Net. The results of these experiments are summarized in Table IX. Notably, ResNet18 has the highest $mAP_{07}$ metric and ResNet50 has the highest $mAP_{12}$ metric.